Determine the cofficent of X5Y10Z5W5 in (x-7y+3z-25)25

1
Expert's answer
2022-02-07T17:16:59-0500

Since the term "(x - 7y +3z - 25)^{25}" do not contain "w", the coefficient of "x^{5}y^{10}z^{5}w^{5}" must be zero.


So we consider "(x - 7y +3z - 2w)^{25}" instead of the original problem "(x - 7y +3z - 25)^{25}" (assuming it a typo error).


We know that by the Multinomial theorem,

"(x - 7y +3z - 2w)^{25} = \displaystyle\sum_{r_1 + r_2 + r_3 + r_4=25}\dfrac{25!}{r_1!r_2!r_3!r_4!}(x)^{r_1}(-7y)^{r_{2}}(3z)^{r_{3}}(-2w)^{r_{4}}"


The coefficient of "x^{5} y^{10}z^{5}w^{5}" in "(x - 7y +3z - 2w)^{25}" is the term obtained by taking "r_1 = 5, r_2 = 10, r_3 = 5, r_4 = 5" in the above summation.


Hence, the coefficient of "x^{5}y^{10}z^{5}w^{5}"

"\begin{aligned}
&= \dfrac{25!}{5!10!5!5!}(1)^{5}(-7)^{10}(3)^{5}(-2)^{5}\\
&=-\dfrac{25!}{5!10!5!5!}(7)^{10}(3)^{5}(2)^{5}\\
\end{aligned}"